OpenStudy (anonymous):

Solve \[\sqrt{3x+1}\] = x-1

OpenStudy (anonymous):

\(\Large \sqrt{3x+1}=(x-1) \) square both sides \(\Large (\sqrt{3x+1})^2=(x-1)^2 \) \(\Large 3x+1=x^2-2x+1 \) can you solve this quadratic from here?

OpenStudy (anonymous):

and especially with radical equations, always check your solution(s) for extraneous solutions.

OpenStudy (anonymous):

Thanks, figured it out. :)